A car moving with constant acceleration covers the distance between two points 60 m apart in 6 sec. It speed as its pass the second point is 15 m/s. what is its speed at the first point?


1
Expert's answer
2021-07-22T10:02:33-0400

By definition the acceleration is given as follows:


"a = \dfrac{v_f - v_i}{t}"

where "v_f = 15m\/s" is the final speed, "v_i" is the initial speed, and "t = 6s" is the time.

The distance travelled under the constant acceleration is given as follows:


"d = v_it + \dfrac{at^2}{2}"

Substituting the expression for "a", have:


"d = v_it + \dfrac{(v_f - v_i)t}{2} = \dfrac{(v_i+v_f)t}{2}\\"

Expressing "v_i", get:


"v_i = \dfrac{2d}{t} - v_f = \dfrac{2\cdot 60m}{6s} - 15m/s = 5m/s"

Answer. 5 m/s.